# Quantitative polynomial cohomology and applications to $\textrm{L}^{p}$-measure equivalence

### v1

@@ Introduction
Measure equivalence (ME) was first introduced by Gromov as a measurable counterpart to quasi-isometry (QI) [gromov, 0.5.E]. It simultaneously generalizes the situation of two lattices sitting inside a locally compact second countable group and that of orbit equivalence (OE). Some standard references include [furman2011survey, gaboriau-survey]. For precise definitions, see Section 1.
@@
Conditions of this kind were first considered by Shalom, first in the form of $\mathrm{L}^{2}$ and $\textrm{L}^{p}$-integrability of lattices [shalom2000rigidity-semisimple] and later in the form of uniform measure equivalence [shalom2004harmonic]. Bader, Furman and Sauer properly defined $\textrm{L}^{p}$-measure equivalence ($\textrm{L}^{p}$-ME, see Section 1 for a definition) in [bader-furman-sauer], which has been more systematically studied and further generalized since [DKLMT].
One can see $\textrm{L}^{p}$-measure equivalence as a way of interpolating between measure equivalence and quasi-isometry (at least among amenable groups). Indeed, Shalom showed that two finitely generated amenable groups are quasi-isometric if and only if they are mutually coubounded $\mathrm{L}^{\infty}$-ME [shalom2004harmonic, 2.1.7]. Mutual coboundedness should be thought of as a fundamental domain compatibility condition that extends OE (see Section 1 for a definition).
A new goal would be to study infinite countable groups up to $\textrm{L}^{p}$-ME. An important result in this direction is due to Bowen, who showed that if two finitely generated groups are $\mathrm{L}^{1}$-ME, then they have the same asymptotic growth [austin2016, Appendix B]. This already shows the diversity of $\mathrm{L}^{1}$-ME classes among amenable groups, as it distinguishes virtually nilpotent groups, groups of intermediate growth, and amenable groups of exponential growth.
@@
One class of groups for which we can give more precise statements about their $\textrm{L}^{p}$-ME classification are (virtually) nilpotent groups. Indeed, Austin showed that finitely generated groups which are $\textrm{L}^{1}$-ME have bilipschitz asymptotic cones [austin2016, 1.1]. Moreover, Pansu showed that all asymptotic cones of a simply connected nilpotent Lie group are isomorphic to its associated Carnot group [pansu89-carnot]. Combining these two results, we obtain that two finitely generated nilpotent groups which are $\textrm{L}^{1}$-ME have the same associated Carnot group. On the other hand, Delabie, Llosa Isenrich and Tessera showed a (slightly stronger) converse to this result: two finitely generated virtually nilpotent groups with isomorphic associated Carnot groups are $\textrm{L}^{p}$-OE for some $p>1$ [delabie-llosa-tessera]. We sum up these results in the following theorem.
@@
[delabie-llosa-tessera, 1.5] Let $\Gamma$ and $\Lambda$ be two finitely generated virtually nilpotent groups. The following are equivalent:
@@
Given two nilpotent groups with the same associated Carnot, one may ask what is the largest $p>1$ for which they are $\textrm{L}^{p}$-OE or $\textrm{L}^{p}$-ME. A first obstruction to $\textrm{L}^{p}$-ME for large $p>1$ was obtained using a particular type of central extensions [delabie-llosa-tessera, 1.13]. Our first result gives an obstruction to mutually cobounded $\textrm{L}^{p}$-ME (and in particular $\textrm{L}^{p}$-OE) when the Betti numbers of the groups in question are different.
@@
The number $p$ can be described explicitly in terms of some data on the groups; we will make this more precise later in this introduction (see Theorem D). For instance, we may consider the simply connected nilpotent Lie group $L_{n}$ whose Lie algebra is given by
@@
$\bullet$ (Furman) Property $(T)$ is invariant under ME [furman1999-ME, 1.4].
$\bullet$ (Gaboriau) $\ell^{2}$-Betti numbers of countable groups are invariant under ME (up to proportionality on the sequence) [gaboriau-invariantsl2, 6.3].
$\bullet$ (Shalom) Betti numbers of finitely generated nilpotent groups are invariant under quasi-isometry [shalom2004harmonic, 1.2] (where quasi-isometry should be understood as mutually cobounded $\mathrm{L}^{\infty}$-ME, so we can see that Theorem A is in fact a strengthening of this result).
$\bullet$ (Sauer) The real cohomology ring of a finitely generated nilpotent groups is invariant under quasi-isometry [sauer-QI-invariants, 1.5].
$\bullet$ (Marrakchi-de la Salle) For $2<p<\infty$, the fixed point property $F\textrm{L}^{p}$ is invariant under $\textrm{L}^{p}$-ME [marrakchi-dlSalle, 7.4].
$\bullet$ (Das) For $1\leq p<\infty$, vanishing of $\ell^{p}$-cohomology in degree 1 is invariant under $\textrm{L}^{p}$-ME [das-conformal, 1.1].
The general strategy of all of these results (including Shalom) is to show a version of Shapiro’s lemma for measure equivalence. More precisely, given two ME groups $\Lambda$ and $\Gamma$ and some isometric representation of $\Lambda$ on a Banach space $V$, we can use the $\Gamma$-action on the $\Lambda$-fundamental domain $X_{\Lambda}$ of the ME-coupling to obtain an isometric representation $I\rho$ of $\Gamma$, which we call the induced representation of $\rho$. There are many possible choices for the space on which the induced representation can act: to remain in the category of isometric representations on Banach spaces, we will always choose $L^{p}(X_{\Lambda},V)$ for some $1\leq p<\infty$. The classical Shapiro lemma [brown-cohomology, III.6.2] states that if $\Lambda$ is a finite index subgroup in $\Gamma$, then for every $\Lambda$-module $V$ there is an isomorphism
@@
In more general settings, in order to ensure that this induction map $I$ with values in $L^{p}(X_{\Lambda},V)$ is well defined at the level of cochains, some quantitative control is required. This control can come either from conditions on the coupling space (that is, sufficient integrability) or from imposing growth restrictions on the cochains. For instance, what Shalom really proved in [shalom2004harmonic] to obtain invariance of Betti numbers of nilpotent groups, is that if $\Lambda$ and $\Gamma$ are two mutually cobounded $\mathrm{L}^{\infty}$-ME groups, then for every unitary representation $(\rho,V)$ of $\Lambda$ we have continuous injective linear maps:
@@
This result should be compared with the result by Monod and Shalom on bounded cohomology [monod-shalom, 4.4], stating that if $\Lambda$ and $\Gamma$ are ME groups, then for every unitary representation $(\rho,V)$ of $\Lambda$ we have that:
$H^{2}_{b}(\Lambda,\rho)\xhookrightarrow{I}H^{2}_{b}(\Gamma,I\rho).$
Here no integrability condition is required on the ME, yet a somewhat similar result is obtained. The reason is that here, the quantitative control is placed on the cochains, which are bounded, and not on the integrability of the coupling. Another result of a similar nature, is the one showed by Marrakchi and de la Salle [marrakchi-dlSalle, 7.4], stating that if $\Lambda$ and $\Gamma$ are finitely generated $\textrm{L}^{p}$-ME groups, then for every isometric representation $(\rho,V)$ on some uniformly convex Banach space $V$, the $\textrm{L}^{p}$-induction module $(I^{p}\rho,I^{p}(V))$ satisfies:
$H^{1}(\Lambda,\rho)\xhookrightarrow{I}H^{1}(\Gamma,I^{p}\rho)$
@@
These ideas were implicitly used by Bader and Sauer in [bader-sauer], who showed that by restricting to cochains that grow at most at polynomial speed and requiring $\textrm{L}^{p}$-integrability for every $p<\infty$, one obtains injectivity of the induction map from a lattice to its ambient locally compact group in [bader-sauer, 6.23].
In order to prove Theorem A, we show that the induction map $I$ is well-defined in the same way as Bader-Sauer, but for a quantitative variant of polynomial cohomology (where we keep track of the degrees of the polynomials) and in the more general setting of ME.
@@
Our main statement regarding induction of polynomial cohomology is the following (which can be stated in more general settings, but under more restrictive and complicated conditions, see Theorem 3.4).
@@
Let $1<p,q<\infty$ such that $\frac{1}{p}+\frac{1}{q}=1$, $k\in\mathbb{N}_{\geq 2}$ and $M,N\in\mathbb{R}_{>0}$, with $M\geq N$. Let $\Lambda$ and $\Gamma$ be two finitely generated virtually nilpotent groups with ME coupling $(\Omega,m,X_{\Lambda},X_{\Gamma})$ and let $d_{\Gamma}=\inf\{d\in\mathbb{R}_{>0},\#S_{\Gamma}(n)\lesssim n^{d}\}<\infty$. Suppose that:
$\bullet$ $X_{\Gamma}\subseteq X_{\Lambda}$,
$\bullet$ the ME coupling is $L^{pM}$-integrable on $\Lambda$,
$\bullet$ the ME coupling is $L^{r}$-integrable on $\Gamma$ for some $r>\frac{d_{\Gamma}kq}{p}+(M+1)q-1$.
Then the induction map $I:C_{\mathrm{pol}\leq M}^{j}(\Lambda,V)^{\Lambda}\to C_{\mathrm{pol}\leq M}^{j}(\Gamma,L^{p}(X_{\Lambda},V))^{\Gamma}$ is well-defined for $j\geq 0$ and there exist maps $T:C^{j}_{\mathrm{pol}\leq M}(\Gamma,\textrm{L}^{p}(X_{\Lambda},V))^{\Gamma}\to C^{j}(\Lambda,V)^{\Lambda}$ for $j\leq k$ such that:
$\bullet$ Both $I$ and $T$ are linear, continuous and commute with differentials.
$\bullet$ They induce continuous maps in cohomology and reduced cohomology, such that the compositions
$H_{\mathrm{pol}:M\to N}^{k}(\Lambda,\rho)\xrightarrow{I}H^{k}_{\mathrm{pol}:M\to N}(\Gamma,I^{p}(\rho))\xrightarrow{T}H^{k}(\Lambda,\rho),$
$\overline{H}_{\mathrm{pol}:M\to N}^{k}(\Lambda,\rho)\xrightarrow{I}\overline{H}^{k}_{\mathrm{pol}:M\to N}(\Gamma,I^{p}(\rho))\xrightarrow{T}\overline{H}^{k}(\Lambda,\rho).$
are just the comparison maps induced by the inclusion $C_{\mathrm{pol}\leq M}^{*}(\Lambda,V)\xhookrightarrow{}C^{*}(\Lambda,V)$.
The main technical point in the proof of Theorem B is showing that the so-called transfer operator $T$, as defined in the proof of [shalom2004harmonic, 3.2.1], is well-defined in our setting.
@@
The next ingredient for proving Theorem A is understanding when the comparison map appearing in Theorem B is an isomorphism. This is exactly the content of the work of Bader and Sauer (originally due to Ji and Ramsey [ji-ramsey]): a comparison theorem between polynomial cohomology and ordinary cohomology [bader-sauer, 6.14]. Assuming a polynomial control on combinatorial versions of the filling functions $\mathrm{cFV}_{\Gamma}^{j}$ of a group $\Gamma$ (for precise definitions, see Section 1), we establish a comparison theorem between quantitative polynomial cohomology and usual cohomology, depending on the best polynomial bounds that the functions $\mathrm{cFV}_{\Gamma}^{j}$ satisfy. We denote:
@@
The proof of Theorem C, just as in the work of Bader-Sauer, consists in revisiting the classical proof showing that group cohomology can be computed using equivariant cohomology of a simplicial complex on which $\Gamma$ acts geometrically, but this time we keep track of the degrees of the polynomials each time we take a filling. Homological algebra is technically more demanding this time as keeping track of the degrees of the polynomials forces us to see spaces as modules over a different algebra each time the degree of the polynomial increases.
Putting Theorems B and C together we can prove Theorem A (with the same strategy as in [shalom2004harmonic, 4.1.1]) and keep track of the constant $p=p(\Gamma,\Lambda,k)$.
@@
(Quantitative version of Theorem A) Let $\Lambda$ and $\Gamma$ be finitely generated virtually nilpotent groups. Let $d_{\Gamma}=\inf\{d\in\mathbb{R}_{>0},\#S_{\Gamma}(n)\lesssim n^{d}\}$ and $k\in\mathbb{N}_{\geq 2}$. Let $N_{\Lambda,k}=\prod_{i=2}^{k}\deg\mathrm{cFV_{\Lambda}^{i}},N_{\Gamma,k}=\prod_{i=2}^{k}\deg\mathrm{cFV_{\Gamma}^{i}}$ and $N=\max\{N_{\Lambda,k},N_{\Gamma,k}\}$ and let $M>\alpha_{k-1}(N)$ where the functions $\alpha_{i}$ are defined in Theorem C.
Suppose that $\Lambda$ and $\Gamma$ are mutually coboundedly $\textrm{L}^{p}$-ME for some
$p>d_{\Gamma}k+2M+1.$
@@
It is worth noting that Theorem C is more versatile and can give results in other contexts. For example, we can complete some of the results by Bader-Sauer on cohomology of non-cocompact lattices in simple Lie groups. Since their work uses polynomial cohomology without controlling the degrees of its polynomials, they can only deal with lattices in higher-rank semisimple groups as these are $\textrm{L}^{p}$-integrable for every $p<\infty$. A non-cocompact lattice $\Lambda$ in a rank 1 simple Lie group $G$ is only $\textrm{L}^{p}$-integrable for $p<\mathrm{Confdim}(\partial G)$ [shalom2000rigidity-semisimple]. Combining this integrability condition, with estimates by Leuzinger [leuzinger-rank1] and Gruber [gruber-filling] on the filling functions of such lattices, we obtain statements similar to those of Bader-Sauer.
@@
This statement implies for instance that vanishing results on the $\textrm{L}^{p}$-cohomology of pinched negatively curved manifolds obtained by Pansu [pansu08] still hold for non-cocompact lattices in rank 1 simple Lie groups.
@@
$\bullet$ $\mathbb{K}=\mathbb{H}$, $1\leq k\leq n-1$ and $1\leq p<\frac{4n+k-1}{2k}$.
We remark that to prove these results we only use the surjectivity statement of Theorem C (whose proof is much easier than that of the isomorphism statement).

### v2

@@ Introduction
Measure equivalence (ME) was first introduced by Gromov as a measurable counterpart to quasi-isometry (QI) [26, 0.5.E]. It simultaneously generalizes the situation of two lattices sitting inside a locally compact second countable group and that of orbit equivalence (OE). Some standard references include [20, 22]. For precise definitions, see Section 1.
@@
Conditions of this kind were first considered by Shalom, first in the form of $\mathrm{L}^{2}$ and $\textrm{L}^{p}$-integrability of lattices [41] and later in the form of uniform measure equivalence [42]. Bader, Furman and Sauer properly defined $\textrm{L}^{p}$-measure equivalence ($\textrm{L}^{p}$-ME, see Section 1 for a definition) in [4], which has been more systematically studied and further generalized since then [15].
One can see $\textrm{L}^{p}$-measure equivalence as a way of interpolating between measure equivalence and quasi-isometry (at least among amenable groups). Indeed, Shalom showed that two finitely generated amenable groups are quasi-isometric if and only if they are mutually cobounded $\mathrm{L}^{\infty}$-ME [42, 2.1.7]. Mutual coboundedness should be thought of as a fundamental domain compatibility condition that extends OE (see Section 1 for a definition).
A new goal would be to study infinite countable groups up to $\textrm{L}^{p}$-ME. An important result in this direction is due to Bowen, who showed that if two finitely generated groups are $\mathrm{L}^{1}$-ME, then they have the same asymptotic growth [3, Appendix B]. This already shows the diversity of $\mathrm{L}^{1}$-ME classes among amenable groups, as it distinguishes virtually nilpotent groups, groups of intermediate growth, and amenable groups of exponential growth.
@@
One class of groups for which we can give more precise statements about their $\textrm{L}^{p}$-ME classification are (virtually) nilpotent groups. Indeed, Austin showed that finitely generated groups which are $\textrm{L}^{1}$-ME have bilipschitz asymptotic cones [3, 1.1]. Moreover, Pansu showed that all asymptotic cones of a simply connected nilpotent Lie group are isomorphic to its associated Carnot group [36]. Combining these two results, we obtain that two finitely generated nilpotent groups which are $\textrm{L}^{1}$-ME have the same associated Carnot group. On the other hand, Delabie, Llosa Isenrich and Tessera showed a (slightly stronger) converse to this result: two finitely generated virtually nilpotent groups with isomorphic associated Carnot groups are $\textrm{L}^{p}$-OE for some $p>1$ [17]. We sum up these results in the following theorem.
@@
[17, 1.5] Let $\Gamma$ and $\Lambda$ be two finitely generated virtually nilpotent groups. The following are equivalent:
@@
Given two nilpotent groups with the same associated Carnot, one may ask what is the largest $p>1$ for which they are $\textrm{L}^{p}$-OE or $\textrm{L}^{p}$-ME. A first obstruction to $\textrm{L}^{p}$-ME for large $p>1$ was obtained using a particular type of central extensions [17, 1.13]. Our first result gives an obstruction to mutually cobounded $\textrm{L}^{p}$-ME (and in particular $\textrm{L}^{p}$-OE) when the Betti numbers of the groups in question are different.
@@
The number $p$ obtained in this Theorem is always >1>1, so in particular the groups $\Lambda$ and $\Gamma$ in question automatically have the same associated Carnot group. Moreover, this number can be described explicitly in terms of geometric/algebraic data of the groups; we will make this more precise later in this introduction (see Theorem D). For instance, we may consider the simply connected nilpotent Lie group $L_{n}$ whose Lie algebra is given by
@@
$\bullet$ (Furman) Property $(T)$ is invariant under ME [19, 1.4].
$\bullet$ (Gaboriau) $\ell^{2}$-Betti numbers of countable groups are invariant under ME (up to proportionality on the sequence) [21, 6.3].
$\bullet$ (Shalom) Betti numbers of finitely generated nilpotent groups are invariant under quasi-isometry [42, 1.2] (where quasi-isometry should be understood as mutually cobounded $\mathrm{L}^{\infty}$-ME, so we can see that Theorem A is in fact a strengthening of this result).
$\bullet$ (Sauer) The real cohomology ring of a finitely generated nilpotent group is invariant under quasi-isometry [40, 1.5].
$\bullet$ (Monod-Shalom) Vanishing of $H_{b}^{2}(\Gamma,\pi)$ for every unitary representation $\pi$ is invariant under ME. Vanishing of $H^{2}_{b}(\Gamma,\ell^{2}(\Gamma))$ is invariant under ME [35, 7.6].
$\bullet$ (Marrakchi-de la Salle) For $2<p<\infty$, the fixed point property $F\textrm{L}^{p}$ is invariant under $\textrm{L}^{p}$-ME [34, 7.4].
$\bullet$ (Das) For $1\leq p<\infty$, vanishing of $\ell^{p}$-cohomology in degree 1 is invariant under $\textrm{L}^{p}$-ME [14, 1.1].
The general strategy of all of these results (including Shalom) is to show a version of Shapiro’s lemma for measure equivalence. More precisely, given two ME groups $\Lambda$ and $\Gamma$ and some isometric representation of $\Lambda$ on a Banach space $V$, we can use the $\Gamma$-action on the $\Lambda$-fundamental domain $X_{\Lambda}$ of the ME-coupling to obtain an isometric representation $I\rho$ of $\Gamma$, which we call the induced representation of $\rho$. There are many possible choices for the space on which the induced representation can act: to remain in the category of isometric representations on Banach spaces, we will always choose $L^{p}(X_{\Lambda},V)$ for some $1\leq p<\infty$. The classical Shapiro lemma [13, III.6.2] states that if $\Lambda$ is a finite index subgroup of $\Gamma$, then for every $\Lambda$-module $V$ there is an isomorphism
@@
In more general settings, in order to ensure that this induction map $I$ with values in $L^{p}(X_{\Lambda},V)$ is well defined at the level of cochains, some quantitative control is required. This control can come either from conditions on the coupling space (that is, sufficient integrability) or from imposing growth restrictions on the cochains. For instance, what Shalom really proved in [42] to obtain invariance of Betti numbers of nilpotent groups, is that if $\Lambda$ and $\Gamma$ are two mutually cobounded $\mathrm{L}^{\infty}$-ME groups, then for every unitary representation $(\rho,V)$ of $\Lambda$ we have continuous injective linear maps:
@@
This result should be compared with the result by Monod and Shalom on bounded cohomology [35, 4.4], stating that if $\Lambda$ and $\Gamma$ are ME groups, then for every unitary representation $(\rho,V)$ of $\Lambda$ we have that:
$H^{2}_{b}(\Lambda,\rho)\xhookrightarrow[]{I}H^{2}_{b}(\Gamma,I\rho).$
Here no integrability condition is required on the ME, yet a somewhat similar result is obtained. The reason is that here, the quantitative control is placed on the cochains, which are bounded, and not on the integrability of the coupling. Another result of a similar nature is the one shown by Marrakchi and de la Salle [34, 7.4], stating that if $\Lambda$ and $\Gamma$ are finitely generated $\textrm{L}^{p}$-ME groups, then for every isometric representation $(\rho,V)$ on some uniformly convex Banach space $V$, the $\textrm{L}^{p}$-induction module $(I^{p}\rho,I^{p}(V))$ satisfies:
$H^{1}(\Lambda,\rho)\xhookrightarrow[]{I}H^{1}(\Gamma,I^{p}\rho)$
@@
These ideas were implicitly used by Bader and Sauer in [5], who showed that by restricting to cochains that grow at most at polynomial speed and requiring $\textrm{L}^{p}$-integrability for every $p<\infty$, one obtains injectivity of the induction map from a lattice to its ambient locally compact group in [5, 6.23].
In order to prove Theorem A, we show that the induction map $I$ is well-defined in the same way as Bader and Sauer, but for a quantitative variant of polynomial cohomology (where we keep track of the degrees of the polynomials) and in the more general setting of ME.
@@
Our main statement regarding induction of polynomial cohomology is the following (which can be stated in more general settings, but under more restrictive and complicated conditions, see Proposition 3.2 and Theorem 3.4).
@@
Let $k\in\mathbb{N}_{\geq 2}$ and $M,N\in\mathbb{R}_{>0}$, with $M\geq N$. Let $\Lambda$ and $\Gamma$ be two finitely generated virtually nilpotent groups with ME coupling $(\Omega,m,X_{\Lambda},X_{\Gamma})$ and let $d(\Gamma)$ denote the degree of polynomial growth of $\Gamma$.
  1. (1)
    Induction (Proposition 3.2). Suppose that $(\Omega,m,X_{\Lambda},X_{\Gamma})$ is $L^{2M}$-integrable on $\Lambda$. Then the induction maps $I=I_{j}:C_{\mathrm{pol}\leq M}^{j}(\Lambda,V)^{\Lambda}\to C_{\mathrm{pol}\leq M}^{j}(\Gamma,L^{2}(X_{\Lambda},V))^{\Gamma}$ are well-defined linear continuous maps that commute with differentials for $j\geq 0$.
  2. (2)
    Transfer (Theorem 3.4). On top of that, suppose that $X_{\Gamma}\subseteq X_{\Lambda}$ and that $(\Omega,m,X_{\Lambda},X_{\Gamma})$ is $L^{r}$-integrable on $\Gamma$ for some $r>d(\Gamma)k+2M+1$. Then there exist continuous linear maps $T=T_{j}:C^{j}_{\mathrm{pol}\leq M}(\Gamma,\textrm{L}^{2}(X_{\Lambda},V))^{\Gamma}\to C^{j}(\Lambda,V)^{\Lambda}$ for $j\leq k$ (that we call transfer maps) that commute with differentials and satisfy $T\circ I=\mathrm{Id}$. The maps $I$ and $T$ induce continuous linear maps in cohomology and reduced cohomology, such that the compositions
    $H_{\mathrm{pol}:M\to N}^{k}(\Lambda,\rho)\xrightarrow{I}H^{k}_{\mathrm{pol}:M\to N}(\Gamma,I^{2}(\rho))\xrightarrow{T}H^{k}(\Lambda,\rho),$
    $\overline{H}_{\mathrm{pol}:M\to N}^{k}(\Lambda,\rho)\xrightarrow{I}\overline{H}^{k}_{\mathrm{pol}:M\to N}(\Gamma,I^{2}(\rho))\xrightarrow{T}\overline{H}^{k}(\Lambda,\rho).$
    are the comparison maps induced by the inclusion $C_{\mathrm{pol}\leq M}^{*}(\Lambda,V)\xhookrightarrow{}C^{*}(\Lambda,V)$.
The main technical point in the proof of Theorem B is item $(2)$, and in particular, it is to show that the transfer maps $T$, defined in the same way as in [42, 3.2.1], are well-defined in our setting.
Recall that the number $d(\Gamma)$ in Theorem B can be computed explicitly from the central series $C^{j}(\Gamma)=[\Gamma,C^{j-1}(\Gamma)]$ of $\Gamma$ using the Bass-Guivarc’h formula
$d(\Gamma)=\sum_{j=1}^{s}j\,\mathrm{rk}(C^{j-1}(\Gamma)/C^{j}(\Gamma)).$
Since Theorem B is mostly interesting for nilpotent groups $\Lambda$ and $\Gamma$ with the same associated Carnot group, they are at least $L^{1}$-ME, so in particular $d(\Lambda)=d(\Gamma)$ (combining [17, 1.5] and [3, Appendix B]).
@@
The next ingredient for proving Theorem A is understanding when the comparison map appearing in Theorem B is an isomorphism. This is exactly the content of the work of Bader and Sauer (originally due to Ji and Ramsey [29]): a comparison theorem between polynomial cohomology and ordinary cohomology [5, 6.14]. Assuming a polynomial control on combinatorial versions of the filling functions $\mathrm{cFV}_{\Gamma}^{j}$ of a group $\Gamma$ (for precise definitions, see Section 1), we establish a comparison theorem between quantitative polynomial cohomology and usual cohomology, depending on the best polynomial bounds that the functions $\mathrm{cFV}_{\Gamma}^{j}$ satisfy. We denote:
@@
The proof of Theorem C, just as in the work of Bader and Sauer, consists in revisiting the classical proof showing that group cohomology can be computed using equivariant cohomology of a simplicial complex on which $\Gamma$ acts geometrically, but this time we keep track of the degrees of the polynomials each time we take a filling. Homological algebra is technically more demanding this time, as keeping track of the degrees of the polynomials forces us to see spaces as modules over a different algebra each time the degree of the polynomial increases.
Putting Theorems B and C together we can prove Theorem A (with the same strategy as in [42, 4.1.1]) and keep track of the constant $p=p(\Gamma,\Lambda,k)$.
@@
(Quantitative version of Theorem A) Let $\Lambda$ and $\Gamma$ be finitely generated virtually nilpotent groups with the same associated Carnot group. Let $d$ be the common degree of polynomial growth of $\Lambda$ and $\Gamma$ and $k\in\mathbb{N}_{\geq 2}$.
Let $N_{k}=\max\{\prod_{i=2}^{k}\deg\mathrm{cFV_{\Lambda}^{i}},\prod_{i=2}^{k}\deg\mathrm{cFV_{\Gamma}^{i}}\}$.
$\bullet$ Suppose that $\Lambda$ and $\Gamma$ admit a mutually cobounded $\textrm{L}^{p}$-ME coupling for some
$p>2d+2N_{2}+3.$
Then $\dim_{\mathbb{R}}H^{2}(\Lambda,\mathbb{R})=\dim_{\mathbb{R}}H^{2}(\Gamma,\mathbb{R})$.
$\bullet$ Suppose that $\Lambda$ and $\Gamma$ admit a mutually cobounded $\textrm{L}^{p}$-ME coupling for some
$p>kd+2(N_{k}^{2}+(k-2)N_{k})+1.$
@@
It is worth noting that Theorem C is more versatile and can give results in other contexts. For example, we can complete some of the results obtained by Bader and Sauer on cohomology of non-cocompact lattices in simple Lie groups [5]. Since their work uses polynomial cohomology without controlling the degrees of its polynomials, they can only deal with lattices in higher-rank semisimple groups as these are $\textrm{L}^{p}$-integrable for every $p<\infty$. A non-cocompact lattice $\Lambda$ in a rank 1 simple Lie group $G$ is only $\textrm{L}^{p}$-integrable for $p<\mathrm{Confdim}(\partial G)$ [41]. We prove the following theorem, which combined with upper bounds for filling functions of lattices in rank 1 simple Lie groups, can be used to obtain statements that are similar to those of Bader and Sauer.
@@
For instance, Theorem E combined with Wenger’s upper bounds for filling functions of Carnot groups [46, 7.3] imply that non-cocompact lattices in $G=\mathrm{Isom}_{0}(\mathbb{H}^{2}_{\mathbb{O}})$ inherit property $T_{3}] [$5, Definition 1.1] from the group $G$.
@@
(Corollary 5.5 in the text) Let $\Lambda$ be a non-cocompact lattice in $G=\mathrm{Isom}_{0}(\mathbb{H}^{2}_{\mathbb{O}})$. Then $\Lambda$ has property $T_{3}]$, that is, for every unitary representation $\pi$ of $\Lambda$, we have $H^{k}(\Lambda,\pi)=0$ for $k=1,2,3$.
This Corollary allows to enlarge the construction of simple Kazhdan groups given in [18] by allowing to take non-uniform lattices as a starting point.
Theorem E combined with sharper estimates by Leuzinger [32] and Gruber [27] for filling functions of rank 1 lattices imply that Pansu’s vanishing results of $\textrm{L}^{p}$-cohomology of pinched negatively curved manifolds [38] still hold for non-cocompact lattices in rank 1 simple Lie groups.
Corollary 0.4.
(Corollary 5.6 in the text) Let $\Lambda$ be a non-cocompact lattice in $G=\mathrm{Isom}_{0}(\mathbb{H}^{n}_{\mathbb{K}})$, where $\mathbb{H}^{n}_{\mathbb{K}}$ denotes the hyperbolic space of dimension $n$ over $\mathbb{K}=\mathbb{R},\mathbb{C}$ or the quaternions $\mathbb{H}$. Let $k\geq 1$ and $1\leq p<\infty$. We have $H^{k}(\Lambda,\ell^{p}(\Lambda))=0$ if
@@
$\bullet$ $\mathbb{K}=\mathbb{H}$, $1\leq k\leq n-1$ and $1\leq p<\frac{4n+k-1}{2k}$,
$\bullet$ $\mathbb{K}=\mathbb{O}$, $k=1,2,3$ and $1\leq p<\frac{15+k}{2k}$.
We remark that to prove Theorem E and its consequences we only use the surjectivity statement of Theorem C (whose proof is much easier than that of the isomorphism statement).